Metabolizing a candy bar releases about 10

6

106 joules of energy. Assume the candy bar's mass is 0.12 kg

kgand use the formula kinetic energy =1

2


mv

2

=12mv2, where m

m is the object's mass and v

v is its velocity. How fast must the candy bar travel to have the same 10

6

106 joules in the form of kinetic energy?


1
Expert's answer
2021-10-05T10:07:05-0400

Rearranging the equation "E=\frac{1}{2}mv^2" to solve for "v", we get:

"v=\sqrt{\frac{2E}{m}}=\sqrt{\frac{2\times10^6J}{0.12kg}}=4.1\times10^3\space{m/s}"


Answer: "4.1\times10^3\space{m/s}"